Find the centre, the foci, and the asymptotes of the hyperbola 4x2 - 9y2 -16x - 54y = 101.
Answer #1

Given hyperbola 40²_gy ²-160-s4y=10) 42²-lon - gy²-suy=101 un uo] -g[y² + 6y] =101 un untu-y]-9 [g² by +9 -9] =10) | u 6-24]Eruction for right-lest Compare with stanelere facing (ho-ni - (1978) 92 with rentre (hik) Hen h=2 K=-3 a=3 b=2 facing parabo